The humidity meter shows the relative humidity of the room air at 78% when the temperature
is 22 ° C. What is the absolute humidity in g / m3 of air and what is Dew point (temperature)?
1
Expert's answer
2019-03-11T12:40:25-0400

1) Calculate the absolute humidity in "\text{ g/m}^3" substituting the relative humidity in % and the temperature in Celsius degrees:


"AH=\frac{6.112\cdot 2.1674\cdot e^{17.67T/(243.5+T)}\cdot RH}{273.15+T}=15.14\text{ g/m}^3."

2) Calculate the parameter "\gamma (T, RH)" for "T=22^\circ\text{C}" (Celsius) and "RH=0.78" (not %) using Magnus formula:


"\gamma (T, RH)=\frac{17.27\cdot T}{237.7+T}+\text{ln}RH=1.215."

Calculate the dew point:


"T_{DP}=\frac{237.7\cdot \gamma (T, RH)}{17.27-\gamma (T, RH)}=17.98^\circ\text{C}."

The formula has a maximum error "\pm 0.4^\circ\text{C}."